# Digital Twin-Empowered Task Assignment in Aerial MEC Network: A Resource Coalition Cooperation Approach with Generative Model

### v2

@@ V-A Coalition Formation Rules
1) Pareto optimal rule: Assuming there is a set of coalition participants $v\subset V$, and there are two coalitions $G_{1}=\left\{g_{1}^{1},g_{2}^{1},\ldots g_{k_{1}}^{1},\ldots g_{K_{1}}^{1}\right\}$ and $G_{2}=\left\{g_{1}^{2},g_{2}^{2},\ldots g_{k_{2}}^{2},\ldots g_{K_{2}}^{2}\right\}$ in $v$. If the utility of $UAV_{j}$ in the coalition set $G_{1}$ is better than $G_{2}$, that is, $u_{i}\left(G_{1}\right)\geq u_{i}\left(G_{2}\right),\forall i\in v$, then the Pareto optimal rule is expressed as $G_{1}\triangleright G_{2}$, if and only if there is at least one $UAV_{j}$ takes a strict inequality.
2) Combining rule: Assuming the coalition set $G_{3}=\left\{g_{1},g_{2},\ldots,g_{k_{b}}\right\}$, if there is $G_{4}=\left\{U_{k_{b}=1}^{K_{b}}g_{k_{b}}\right\}\triangleright G_{3}$, $G_{3}$ will be formed into $G_{4}$, that is, $G_{3}\succ G_{4}$.
3) Dividing rule: Assuming the coalition set $G_{5}=\bigcup_{k_{s}=1}^{K_{s}}g_{k_{s}}$, if there is $G_{6}=\left\{g_{1},g_{2},\ldots,g_{k_{s}}\right\}\triangleright G_{5}$, $G_{5}$ will be split into $G_{6}$, that is, $G_{5}\prec G_{6}$.
@@ V-B Resource Coalition Cooperation Algorithm
$n(t)=\left\{\boldsymbol{B}_{V}^{t},\boldsymbol{C}_{i}^{t},DT_{MED_{i}}^{t},\boldsymbol{F}_{V}^{t},\mathcal{C}_{V}^{t},Env^{t}\right\}$ (16)
@@
Step 1. Collected status information of the UAVs: The status information of the multi-UAV is collected by the airship. The DT of the multi-UAV constructs the initial network status $n(t=0)$ based on the collected status information.
Step 2. Initial resource pool and coalition set: The resource scheduler forms each $DT^{t}_{UAV_{j}}$ into a coalition $g_{j}$, so that generates the initial coalition set $G_{0}=\{g_{1},g_{2},\ldots,g_{j},\ldots,g_{n}\}$.
Step 3. Stable coalition structure with maximization of coalition utility: Based on the status information of the task assignment, all coalitions maximize the utility of each coalition participant by combining or dividing. When several coalitions satisfy the rules of combining, these coalitions will be reorganized into a larger coalition, i.e., $G^{{}^{\prime}}_{0}=\{g^{{}^{\prime}}_{1},g^{{}^{\prime}}_{2},g^{{}^{\prime}}_{3},g^{{}^{\prime}}_{j}\}$. When a coalition satisfies the dividing rule, the coalition is split into smaller coalitions.
Step 4. Constructed new coalitions: Resource scheduler selects the coalition with the optimal utility as the input for resource coalition cooperation, and then determines the participants in the coalition. When the current time slot ends, immediately abandon the original coalition construction and return to the Step 2.
V-C DT Deviation Analysis
@@
$\tilde{E}_{V}^{t}=E^{t}_{V,tr}+\sum_{j\in V}\left(p^{t}_{j,cp}{\tilde{T}}^{t}_{j,cp}+p^{t}_{j,h}{\tilde{T}}_{V}^{t}\right)$ (18)
@@
VI Simulation Results and Discussion
This section is to verify the performance advantages of the resource coalition cooperation and task assignment strategy proposed in this paper. We perform numerical simulation and analysis in aspects such as DT fidelity, resource utilization, energy consumption, and coalition utility.
TABLE II: Simulation experiment parameter setting
@@ VI Simulation Results and Discussion
VI-A Simulation Scenario and Parameter Settings
Aim to make the experiment better focused on the resource coalition cooperation strategy proposed in this paper, we select an area of 1 km * 1 km set that is uncovered by an airship and 30 UAVs in the data set. The simulation environment considers all UAVs flying at the fixed altitude of 800 m, an airship hovers at a fixed altitude of 2000 m. The transmission channels of M2U and U2A in this experimental design of an aerial MEC network adopt the Rayleigh fading model. The number of time slots is set to 500 and each slot has a maximum duration $t$ = 60 s. The DTs of UAVs and MEDs are established on the airship. The DTs of MEDs reflect the dynamic demand and supply of resources by the UAVs. We feed about two hundred labeled strategies to the diffusion model and train the denoising process of U-Net with maximum-likelihood estimation [28]. The tasks were generated randomly on the MEDs and assigned to UAVs with airship assistance. The powers of UAV hovering and other parameters are listed in Table II [31].
Moreover, in order to reflect the advantages of proposed method, we compare the performance of the proposed approach with the following alternatives, i.e., the cooperation-based grand coalition method [34] and the noncooperation-based Nash equilibrium method [35] as the benchmarking method. Grand coalition means all UAVs cooperate with each other for task offloading in MED. Nash equilibrium means the UAV independently provides task offloading services for MED, requiring a careful evaluation of its utility in multi-UAV.
@@ VI-A Simulation Scenario and Parameter Settings
VI-B Implementation of One-to-Many Task Assigning
@@ VI-C Effectiveness Analysis of DT and GM
Figure 6: The total number of U2U with different numbers of UAVs.
VI-D Resource Utilization Results and Analysis
@@
VI-E Energy Consumption Results and Analysis
@@
VI-F Utility Results and Analysis
@@ VII Conclusion
In future work, we intend to explore multi-agent reinforcement learning techniques to deepen our understanding of the task assignment strategy between MEDs and UAVs, particularly regarding the dynamics of unknown load levels in the aerial MEC network.

### v3

@@ V-A Coalition Formation Rules
Definition 1: (Pareto optimal rule) Assuming there is a set of coalition participants $v\subset V$, and there are two coalitions $G_{1}=\left\{g_{1}^{1},g_{2}^{1},\ldots g_{k_{1}}^{1},\ldots g_{K_{1}}^{1}\right\}$ and $G_{2}=\left\{g_{1}^{2},g_{2}^{2},\ldots g_{k_{2}}^{2},\ldots g_{K_{2}}^{2}\right\}$ in $v$. If the utility of $UAV_{j}$ in the coalition set $G_{1}$ is better than $G_{2}$, that is, $u_{i}\left(G_{1}\right)\geq u_{i}\left(G_{2}\right),\forall i\in v$, then the Pareto optimal rule is expressed as $G_{1}\triangleright G_{2}$, if and only if there is at least one $UAV_{j}$ takes a strict inequality.
Definition 2: (Combining rule) Assuming the coalition set $G_{3}=\left\{g_{1},g_{2},\ldots,g_{k_{b}}\right\}$, if there is $G_{4}=\left\{U_{k_{b}=1}^{K_{b}}g_{k_{b}}\right\}\triangleright G_{3}$, $G_{3}$ will be formed into $G_{4}$, that is, $G_{3}\succ G_{4}$.
Definition 3: (Dividing rule) Assuming the coalition set $G_{5}=\bigcup_{k_{s}=1}^{K_{s}}g_{k_{s}}$, if there is $G_{6}=\left\{g_{1},g_{2},\ldots,g_{k_{s}}\right\}\triangleright G_{5}$, $G_{5}$ will be split into $G_{6}$, that is, $G_{5}\prec G_{6}$.
@@ V-B Resource Coalition Cooperation Algorithm
$n(t)=\left\{\boldsymbol{B}_{V}^{t},\boldsymbol{C}_{i}^{t},DT_{MED_{i}}^{t},\boldsymbol{F}_{V}^{t},\mathcal{C}_{V}^{t},Env^{t}\right\},$ (17)
@@
Step 1 (Line 1-line 5). Collected status information of the UAVs: The status information of the multi-UAV is collected by the airship. The DT of the multi-UAV constructs the initial network status $n(t=0)$ based on the collected status information.
Step 2 (Line 6). Initial resource pool and coalition set: The resource scheduler forms each $DT^{t}_{UAV_{j}}$ into a coalition $g_{j}$, so that generates the initial coalition set $G_{0}=\{g_{1},g_{2},\ldots,g_{j},\ldots,g_{n}\}$.
Step 3 (Line 7-line 15). Stable coalition structure with maximization of coalition utility: Based on the status information of the task assignment, all coalitions maximize the utility of each coalition participant by combining or dividing. When several coalitions satisfy the rules of combining, these coalitions will be reorganized into a larger coalition, i.e., $G^{{}^{\prime}}_{0}=\{g^{{}^{\prime}}_{1},g^{{}^{\prime}}_{2},g^{{}^{\prime}}_{3},g^{{}^{\prime}}_{j}\}$. When a coalition satisfies the dividing rule, the coalition is split into smaller coalitions.
Step 4 (Line 19). Constructed new coalitions: Resource scheduler selects the coalition with the optimal utility as the input for resource coalition cooperation, and then determines the participants in the coalition. When the current time slot ends, immediately abandon the original coalition construction and return to the Step 2.
VI Performance Analysis
In this section, the stability and convergence of TU coalition game are illustrated. Then, we define the fidelity of DT model and analyse its deviation.
VI-A Stable Coalition Definition and Proof
[…]
Proof: The system model in this paper is a discrete periodic synchronization time slot system, and network status updates only occur at the end of the synchronization time slot. We want to prove that the TU coalition game algorithm proposed in this paper can converge to a unique and ${\mathbb{D}}_{hp}$ stability coalition within a limited number of iterations. The specific proof is as follows
a) If $\forall g_{k}\in G$ is split into $\{g_{1}^{{}^{\prime}},g_{2}^{{}^{\prime}},\ldots,g_{q}^{{}^{\prime}},\ldots,g_{Q}^{{}^{\prime}}\}$ and there is $\{g_{1},g_{2},\ldots,g_{q},\ldots,g_{Q}\}\rhd g_{k}$. Then according to the dividing rule coalition $G\prec G^{{}^{\prime}}=\{g_{1},g_{2},\ldots,g_{1}^{{}^{\prime}},g_{2}^{{}^{\prime}},\ldots,g_{q}^{{}^{\prime}},\ldots,g_{Q}^{{}^{\prime}},\ldots,g_{K}\}$, the split result is inconsistent with the coalition set $G$.
b) If there is $\cup_{q\in Q}g_{q}\rhd\left\{g_{q}\mid q\in Q\right\}$, then according to the coalition combining rules, $G\succ G^{{}^{\prime\prime}}=\left\{\cup_{q\in Q}g_{q}\right\}\cup\left\{g_{q}\mid q\in\{1,2,\ldots,K\}\backslash Q\right\}$ can be obtained. However, the combining result is contradictory to $G$.
The above assumptions a) and b) further illustrate that as long as the coalition is stable, then the coalition must be the only coalition structure in the current network status.
VI-B Convergence Analysis
Suppose the initial coalition is $G_{0}$, and the set of stable coalitions after the $x$th optimization is $G_{x}$. Then, the coalition combining and dividing rules are executed, and $G_{x}$ is formed after the $(x+1)$th combining or dividing of the coalition set $G_{x+1}$ that has not been iteratively traversed. Similarly, the initial coalition $G_{0}$ becomes $G_{X}$ after $X$ rounds of combine or divide iterations. Obviously, for a set $V$ composed of $n$ coalition participants, under the limitation of the objective function of optimizing coalition utility with multiple constraints, the number of coalitions constructed must be enumerable and can be solved using Bell numbers. At the same time, with the optimization iteration of the coalition, the TU coalition game method proposed in this paper can also tend to converge in a limited number of iterations, forming a stable coalition in the current time slot, thereby providing task assignment decisions for MED.
VI-C DT Deviation Analysis
@@
$\tilde{E}_{V}^{t}=E^{t}_{V,tr}+\sum_{j\in V}\left(p^{t}_{j,cp}{\tilde{T}}^{t}_{j,cp}+p^{t}_{j,h}{\tilde{T}}_{V}^{t}\right),$ (19)
@@
VII Simulation Results and Discussion
This section is to verify the performance advantages of the resource coalition cooperation and task assignment strategy proposed in this paper. We perform numerical simulation and analysis in aspects such as effectiveness of DT and GM, resource utilization, energy consumption, and coalition utility.
TABLE III: Simulation experiment parameter setting
@@ VII Simulation Results and Discussion
VII-A Simulation Settings
Aim to make the experiment better focused on the resource coalition cooperation strategy proposed in this paper, we select an area of 1 km * 1 km set that is covered by an airship and multiple UAVs in the data set. The simulation environment is mainly for the needs of densely populated urban high-rise buildings. The existing most urban high-rise buildings are below 800 m. Therefore, we consider all UAVs flying at a fixed altitude of 800 m. Furthermore, in addition to cooperating with the UAVs below, the airship will also carry some observation payloads for air-to-ground monitoring. At this time, its own safety factors must also be considered. For example, in the case of sudden volcanic disasters, the height of the smoke column can reach hundreds or even thousands of meters. In this way, we set an airship to hover at a fixed altitude of 2000 m. The transmission channels of M2U and U2A in this experimental design of an aerial MEC network adopt the Rayleigh fading model. The number of time slots is set to 500 and each slot has a maximum duration $t$ = 60 s. The DTs of UAVs and MEDs are established on the airship. The DTs of MEDs reflect the dynamic demand and supply of resources by the UAVs. Hence, we turn the non-convex resource optimization problem into a convex optimization problem with linear constraints that can be solved using conventional optimization tools such as CVX. Moreover, we feed about two hundred labeled strategies to the diffusion model and train the denoising process of U-Net with maximum-likelihood estimation [39]. The tasks were generated randomly on the MEDs and assigned to UAVs with airship assistance. We perform simulation experiments in the platform of MATLAB R2019a. The PC is equipped with the 64-bit Microsoft Windows 10 operating system, 16 GB RAM, and Intel(R) Core(TM) i5-8250U CPU. To evaluate the efficiency of our proposed solution, the utilization of computing resource (12), the total energy consumption Eq. (7), and the participant utility Eq. (10) of our proposed scheme with these of benchmark schemes are compared under four different settings, i.e., the computational complexity $\vartheta^{t}_{i}$, the number of UAV $n$, and the number of time slots $\tau$. The powers of UAV hovering and other parameters are listed in Table III [42].
Moreover, in order to reflect the advantages of proposed method, we compare the performance of the proposed approach with the following alternatives, i.e., the cooperation-based grand coalition method [19] and the noncooperation-based Nash equilibrium method [11] as the benchmark method.
Grand coalition: Grand coalition means all UAVs cooperate with each other to form a multi-UAV system to provide all-to-one task-executing services. In this way, the coalition enhances the quality of services. However, it can also lead to poor utilization of UAV resources. And this approach does not pay attention to the utilization of all resources.
Nash equilibrium: Nash equilibrium means the UAV independently provides task offloading services for MED, requiring a careful evaluation of its utility in multi-UAV. Each UAV provides task offloading services independently, requiring careful evaluation of the inter-UAV relationships. In other ways, additional information interaction among multiple UAVs is required when generating resource cooperation strategies.
@@ VII-A Simulation Settings
Figure 6: The total number of U2U with different numbers of UAVs.
VII-B Implementation of One-to-Many Task Assigning
@@ VII-C Effectiveness Analysis of DT and GM
VII-D Resource Utilization Results and Analysis
@@
VII-E Energy Consumption Results and Analysis
@@
VII-F Utility Results and Analysis
@@ VIII Conclusion
In future work, we intend to explore the collaboration between multiple heterogeneous UAVs, while the other is the application of generative models in specific tasks, such as assisting the generation of oblique photography images and wireless spectrum maps.